OpenStudy (loser66):

2 functions a, b : N --> N are defined \(a_0 =2; a_1 =2; a_{n+2}=a_{n+1} + a_n\) \(b_0=2, b_1=2, b_{n+2} = b_{n+1} * b_n\) Prove by induction that \(b_n \geq a_n\) Please, help

OpenStudy (loser66):

For n = 0, \(a_2 = a_1 + a_0 = 4\\a_3 = 6 \\a_4 = 10\\a_5= 16\\a_6 = 26\), Hence the sequence is {2,2,6,10,16,26,.....}

OpenStudy (anonymous):

\[b _{2}=4\]\[b _{3}=8\]\[b _{4}=32\]\[b _{5}=256\]

OpenStudy (loser66):

Don't we have to find explicit formula for \(a_n, b_n\)?

OpenStudy (anonymous):

i don't use any

OpenStudy (loser66):

Ok, show me more, please

OpenStudy (loser66):

My attempt: For a_n Characteristic equation: r^2 -r -1 =0, \(r =\dfrac{1\pm\sqrt5}{2}\), Then the solution is \(a_n = A (\dfrac{1+\sqrt5}{2})^n + B(\dfrac{1-\sqrt5}{2})^n\) If n = 0, \(a_0 = 2 \), hence A + B = 2 If n=1, \(a_1 = 2, hence ~~ A (\dfrac{1+\sqrt5}{2}) + B(\dfrac{1-\sqrt5}{2})=2\) We can solve for A and B to get the explicit formula for \(a_n\) Do the same to find \(b_n\) From then, use induction to prove that \(b_n \geq a_n\) Am I right so far?

ganeshie8 (ganeshie8):

solving the recurrence relations looks more complicated than necessary

ganeshie8 (ganeshie8):

First of all, notice that both the given sequences are increasing and each term is \(\ge 2\)

ganeshie8 (ganeshie8):

Base case : \(a_0 = 2\ge 2=b_0\) Induction Hypothesis : assume \(b_k\ge a_k\) Induction step : \( b_{k+1} = b_{k} * b_{k-1}\\~\\ \ge a_k*b_{k-1}\\~\\ \ge a_k*2\\~\\ = a_k + a_k\\~\\ \ge a_k + a_{k-1}\\~\\ =a_{k+1} \)

ganeshie8 (ganeshie8):

that should do

OpenStudy (loser66):

WWWWWWWWWoah!! you make it easy!@!:(:)

ganeshie8 (ganeshie8):

maybe work two terms in the base case because the recurrence relation is two levels deep

OpenStudy (loser66):

Got you, thank you so much.

OpenStudy (fibonaccichick666):

I like ganeshie's answer here.

OpenStudy (loser66):

I have a question on residue. I need find the least residue of 1! +2! +...+10! modulo 2.

OpenStudy (fibonaccichick666):

But I would add a corollary that shows \(a_{n+1}\ge a_n\)

OpenStudy (fibonaccichick666):

or k in this instance

ganeshie8 (ganeshie8):

that would make it more complete...

OpenStudy (loser66):

Can I argue: 2! = 0 mod 2 other are same and 1! =1 mod 2 then 1! +2!+.....+10! = 1 mod 2 Is it ok?

ganeshie8 (ganeshie8):

looks good to me

OpenStudy (loser66):

The problem is I don't find any theorem or corollary in book support my argument.

ganeshie8 (ganeshie8):

which argument ?

OpenStudy (fibonaccichick666):

I would say your residue argument relies solely on the properties of even and odd numbers

ganeshie8 (ganeshie8):

yeah we have, \(2\mid n!\) for all \(n\ge 2\). so \(n!\equiv 0 \pmod{2}\) for all \(n\ge 2\)

OpenStudy (fibonaccichick666):

and definition of factorial, since they must all be even

OpenStudy (loser66):

\(1! \equiv 1(mod2)\\2!\equiv 0 (mod 2)\\--------------------\\1! +2! \equiv 1 (mod 2)\) This property.

OpenStudy (fibonaccichick666):

That's a property? I thought it was just computation

OpenStudy (fibonaccichick666):

since 1+2=3 and 2 is congruent to 1mod2

OpenStudy (fibonaccichick666):

3*

OpenStudy (loser66):

Since I have to find the least residue of that sum modulo 7,12, 25 and I don't want to expand the factorial. I need some theorem allow me to add the left hand side together and add the right hand side together but I don't have it on my book. :(

OpenStudy (loser66):

I have something like \(x \equiv 2 (mod n)\\x+2\equiv 2+2 (mod n)\)

OpenStudy (fibonaccichick666):

Ahh, instate like chinese remainder them?

OpenStudy (fibonaccichick666):

that's the only thing I recall that allows addition/mult on the mod side

OpenStudy (loser66):

It means I can add the same number both sides of the equivalence,

OpenStudy (fibonaccichick666):

shouldn't the least res always be 1?

OpenStudy (fibonaccichick666):

strike that, I can't prove it

ganeshie8 (ganeshie8):

if \(a\equiv c \pmod{n}\) and \(b\equiv d \pmod{n}\), then : \(a+b \equiv c+d \pmod{n}\)

ganeshie8 (ganeshie8):

you want to prove that ?

OpenStudy (loser66):

oH, I have it also. oh, yeah!!

OpenStudy (fibonaccichick666):

I mean, we can prove it is always an odd value for the first one but for the 7,how do we know zero is necessarily not occurring. and for 25 as well

OpenStudy (fibonaccichick666):

since we can essentially write it as 1+2(n+1) where n+1= 2!/2+3!/2+...+10!/2